OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

hero (hero):

1. Isolate x in the first equation 2. Replace the x in the second equation with the expression for x isolated in the first equation.

OpenStudy (anonymous):

its a ?

OpenStudy (anonymous):

you have x=1+4y putting this into 2x-9y=3 gives you 2(1+4y)-9y=3 <=> 2+8y-9y=3 <=> -y=1 <=> y=-1 Now put y into the first equation: x=1+4(-1)=1-4=-3 so you have (-3,-1)

OpenStudy (anonymous):

ty i jus dont understand how u put the first equation into the second?

hero (hero):

1st equation: x - 4y = 1 2nd equation: 2x - 9y = 3 1st equation (x isolated): x = 1 + 4y 2nd equation (x substituted): 2(1 + 4y) - 9 = 3